OpenStudy (anonymous):

Write the slope intercept form of the equation of the line described. Through: (3,-2), parallel to y=-1/3x+4

OpenStudy (anonymous):

Parallel lines have the same slope and from the equation given you can tell the slope is -1/3

OpenStudy (anonymous):

The slope intercept form?

OpenStudy (anonymous):

To get that you use the point you were given plus the slope in this formula y-y1=m(x-x1) y+2=-1/3(x-3) y+2=-1/3x+3 y=-1/3x+1

OpenStudy (anonymous):

Thankx